OpenStudy (anonymous):

Tori invested $8000 in money market funds. Part was invested at 5% simple interest and the rest was invested at 7% simple interest. At the end of the year, Tori had earned $496 in interest. How much did she invest in each fund?

OpenStudy (anonymous):

@aliciabennett We need some variables and algebra to do this problem.

OpenStudy (anonymous):

I will pick \(x\) and \(x\) to be our variables.

OpenStudy (anonymous):

I mean \(x\) and \(y\)

OpenStudy (anonymous):

this is exactly the way it is written in my book

OpenStudy (anonymous):

Yeah, the book is fine.

OpenStudy (anonymous):

I am thinking that at 5% she invested $3200 and at 7% she invested 4800

OpenStudy (anonymous):

Does it work out to the right total?

OpenStudy (anonymous):

Let \(x\) be the amount in the \(5\%\) account and \(y\) be the ammount in the \(7\%\) account.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

"Tori invested $8000 in money market funds." This gives us the equation:\[ x + y = $8000 \]

OpenStudy (anonymous):

Does this equation make sense?

OpenStudy (anonymous):

Okay nevermind. Your solution is correct. How did you get it?

OpenStudy (campbell_st):

well all you need is let x be the 1st investment and then 8000 - x is the other investment using I = Prn \[ 496 = x \times 0.05 \times 1 + (8000 - x)\times 0.07 \times 1\] \[496 = 0.05x + 560 - 0.07x\] so your equation becomes 496 = 560 - 0.02x just solve for x which will tell you how much is invested at 5% and the balance is invested at 7%. hope this helps